OpenStudy (anonymous):

How does one antidifferentiate x*(x^2+1)^(-3/2)? Thanks in advance, Nick

OpenStudy (anonymous):

if i help u will u help me

OpenStudy (anonymous):

try using substitution.

OpenStudy (anonymous):

Sure, as long as it is the right answer :)

OpenStudy (anonymous):

is there anywhere you see a derivative of some function multiplied by the function itself in the problem?

OpenStudy (anonymous):

subsatute every thing and theres ur answer

OpenStudy (anonymous):

Lilith, I haven't learned substitution yet, and I want to try without it

OpenStudy (anonymous):

Although, I will be soon.

OpenStudy (anonymous):

ok. well then you can't antidifferentiate without factoring out the polynomial. so how would you factor out (x^2)^(-3/2) first?

OpenStudy (anonymous):

Well I took (x^2)^3, which was x^6 + 3x^4 + 3x^2 + 1. And then I figured trying to take the square root of that might be beneficial. But I've been similarly stuck on this step to :/

OpenStudy (anonymous):

i took a second look. because we have (x^2+1) as the base we cannot factor without have imaginary numbers

OpenStudy (anonymous):

this means you have to use substitution. i can walk you through it if you would like

OpenStudy (anonymous):

I'll give you a best response, but no thank you :) I'm certain there is a way to do this without substitution. Else this 1980's textbook is defective..

OpenStudy (anonymous):

ok :) i will get back if i can figure a way out without substitution :)

OpenStudy (anonymous):

Sounds good!

zepdrix (zepdrix):

Nick there is a method called "Advanced guessing", it allows you to skip doing a substitution if you're able to recognize the term and it's derivative in your problem. It's based off of intuition and having a foundation of understanding the substitution method. You cannot solve this without first learning the substitution method.

OpenStudy (anonymous):

that is what i had thought.

OpenStudy (anonymous):

you could also do a taylor series expansion to approximate the function but that, too, will most likely be a bit complicated

zepdrix (zepdrix):

Oh boy that would suck XD lol

OpenStudy (anonymous):

If it helps, this is a verification problem stating: The integral from a to b of x*(x^2+1)^(-3/2) = -1*(x^2+1) from a to b.

OpenStudy (anonymous):

And thanks for all the insight so far!

OpenStudy (anonymous):

Woops, it actually says: The integral from a to b of x*(x^2+1)^(-3/2) = -1*(x^2+1)^(-1) from a to b.

OpenStudy (anonymous):

FOR REAL, it says: The integral from a to b of x*(x^2+1)^(-3/2) = -1*(x^2+1)^(-1/2). Hence the negative square root at the end.

zepdrix (zepdrix):

\[\large -(x^2+1)^{-1/2}\]This should be the answer :D It sounds like the book is just providing you with the answer, it doesn't give any steps? But yah, it's based off of letting u=x^2+1 Then when you take the derivative of your substitution and you're able to replace some stuff and you'll have a much easier anti-derivative to take at that point.

zepdrix (zepdrix):

lol darn you beat me to it ^^

OpenStudy (anonymous):

Haha, well, you have credibility! This book is pretty irritating, it's not listing any steps to achieve verification..Maybe I'll cheat and just use substitution for now.

OpenStudy (anonymous):

that's what i'd do.

zepdrix (zepdrix):

Oh that's cheating? You're suppose to solve the problem without using substitution? Or do you say that because you haven't learned it yet? :D Sounds like maybe the teacher just assigned that problem by accident XD heh

OpenStudy (anonymous):

Yeah, haven't learned it yet. Following this near-vintage 80's book, but homeschooled :)

zepdrix (zepdrix):

Oh fun ^^